A freight train leaves a station travelling at 30 mph A passenger train leaves 1 hours later travelling at 50 mph. At what time will the passenger train overtake the freight train?
Answer by checkley77(12844)

30t=50(t-1)
30t=50t-50
30t-50t=-50
-20t=-50
t=-50/-20
t=2.5 hours after the freight train leaves the station they will meet.